OpenStudy (anonymous):

Multiply, using special product methods. ( x-5)^2 Medal will be given.

jimthompson5910 (jim_thompson5910):

hint: \[\Large (a-b)^2 = a^2 - 2ab + b^2\]

OpenStudy (anonymous):

whats next?

jimthompson5910 (jim_thompson5910):

what are a and b in this case?

OpenStudy (anonymous):

a is x and b is -5 ?

jimthompson5910 (jim_thompson5910):

well when talking about (a-b)^2, b is +5, but I see how you got -5

jimthompson5910 (jim_thompson5910):

so... \[\Large (a-b)^2 = a^2 - 2ab + b^2\] \[\Large (x-5)^2 = x^2 - 2*x*5 + 5^2\] \[\Large (x-5)^2 = x^2 - 10x + 25\]

OpenStudy (anonymous):

so you solve the last part or just leave it like that?

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

the 3rd statement